OpenStudy (anonymous):

A company has a machine that can complete a job in 6 hours. With the help of a newer machine, the job can be completed in 2 hours. How long would the new machine take to do the job alone?

OpenStudy (anonymous):

3 hours

OpenStudy (vijay):

3 hours

OpenStudy (anonymous):

positive its 3?

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

would u like an explanation?

OpenStudy (anonymous):

if its quick sure

OpenStudy (anonymous):

well the 1st machine does 1/6th of the work per hour. but u want the machine to work together to do half of the work per hour. so basicallly 1/6 +x=1/2 and x =1/3 so the new machine does 1/3 of the work per hour. thus 3 hours to finish the workby itself

OpenStudy (anonymous):

ah, ok i get it now thanks